OpenStudy (anonymous):

Hello, I would appreciate if someone could help me with this trig problem.. Deals with identities sin θ - sin θ•cos^2 θ = sin^3 θ

Nnesha (nnesha):

\[\huge\rm cos^2 \theta = ??\]

OpenStudy (anonymous):

cos^2 = 1-sin^2 right....

Nnesha (nnesha):

yep! :-) so replace cos^2 by 1- sin^2 :-)

OpenStudy (anonymous):

okay, after that do i just factor a sinx?

Nnesha (nnesha):

\[\huge\rm sin \theta - \sin \theta \times \color{red}{ 1 - \sin^2 \theta }\] multiply 1 - sin^2 by sin

Nnesha (nnesha):

\[\huge\rm sin \theta - \sin \theta \times \color{red}{( 1 - \sin^2 \theta )}\]let's put parentheses

OpenStudy (anonymous):

--> sinx(sinx-sin^3x)=sin^3

Nnesha (nnesha):

hmm where is negative sign ?? who kidnapped NEGATIVE sign !!

OpenStudy (anonymous):

oops sorry bout that..

OpenStudy (anonymous):

what would i do next?

Nnesha (nnesha):

\[\huge\rm sin \theta \color{blue}{-}\color{red}{ [}\sin \theta \times \color{red}{( 1 - \sin^2 \theta )]}\] so sin(1-sin^2) = sin - sin^3 \[\huge\rm sin \theta \color{blue}{-}\color{red}{ [} \color{red}{( sin \theta - sin^3 \theta )]}\] distribute bracket by negative sign

Nnesha (nnesha):

why did you delete that ?? :(

OpenStudy (anonymous):

Because i made a mistake... lol

OpenStudy (anonymous):

add sin^3x to make the right side 2sin^3x

Nnesha (nnesha):

you have to verify both sides should be equal so don't need to touch right side

Nnesha (nnesha):

verify identities **

OpenStudy (anonymous):

oh okay

Nnesha (nnesha):

let me what u get after distributing bracket by negative sign ?

OpenStudy (anonymous):

sinx-sinx+sin^3x=sin^3x

OpenStudy (anonymous):

sinx-sinx= 0

Nnesha (nnesha):

yes right :-)

OpenStudy (anonymous):

leaving sin^3x=sin^3x

Nnesha (nnesha):

gO_Od to go! :-) nice job!

OpenStudy (anonymous):

Thank you so much!

Nnesha (nnesha):

my pleasure :-) btw \(\huge\color{Green}{{\rm welcome}\rm~to~open~study!!!!}\)\(\Huge \color{gold}{\star^{ \star^{\star:)}}}\Huge \color{green}{\star^{ \star^{\star:)}}}\) \(\Huge \color{blue}{\star^{ \star^{\star:)}}}\Huge \color{red}{\star^{ \star^{\star:)}}}\)\(\rm\color{green}{o^\wedge\_^\wedge o}\)

OpenStudy (phi):

for what it's worth , you could do sin θ - sin θ•cos^2 θ = sin^3 θ factor out sin on the left side to get sin θ ( 1- cos^2 θ ) = sin^3 θ use the identity to replace 1- cos^2 with sin^2 sin θ sin^2 θ = sin^3 θ simplify the left side sin^3 θ = sin^3 θ which shows the identity